OpenStudy (anonymous):

What is the equation of the line perpendicular to the line y = x – 2 that passes through the point (4,3)?

OpenStudy (anonymous):

your line has slope 1, the perpenducular line will have slope \(-1\) so it will be \[y-3=-(x-4)\] or \[y-3=-x+4\] or even \[y=-x+7\]

OpenStudy (anonymous):

which one??

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

@satellite73